# Training robust and generalizable quantum models

### v3

@@ Abstract
Adversarial robustness and generalization are both crucial properties of reliable machine learning models. In this paper, we study these properties in the context of quantum machine learning based on Lipschitz bounds. We derive parameter-dependent Lipschitz bounds for quantum models with trainable encoding, showing that the norm of the data encoding has a crucial impact on the robustness against data perturbations. Further, we derive a bound on the generalization error which explicitly involves the parameters of the data encoding. Our theoretical findings give rise to a practical strategy for training robust and generalizable quantum models by regularizing the Lipschitz bound in the cost. Further, we show that, for fixed and non-trainable encodings, as those frequently employed in quantum machine learning, the Lipschitz bound cannot be influenced by tuning the parameters. Thus, trainable encodings are crucial for systematically adapting robustness and generalization during training. The practical implications of our theoretical findings are illustrated with numerical results.
1 Introduction
@@ A Lipschitz bound of a model $f$ is any $L>0$ satisfying
for all $x_{1},x_{2}\in\mathcal{D}\subseteq\mathbb{R}^{d}$. By definition, Lipschitz bounds quantify the worst-case output change that can be caused by data perturbations and, thus, they provide a useful measure of adversarial robustness. Therefore, they are a well-established tool for characterizing robustness and generalization properties of ML models [2, 6, 9, 10, 11, 12, 13, 14, 15]. Lipschitz bounds cannot only be used to better understand these two properties, but they also allow one to improve them by regularizing the Lipschitz bound during training [2, 6, 16, 17, 18].
Figure 1: Schematic illustration of the quantum model and training setup considered in this work for an exemplary Fashion MNIST data set [19]. The data, $x$, enter the quantum circuit via a trainable encoding, i.e., they are encoded into unitary operators $U_{j,\Theta_{j}}(x)$ via an affine function $w_{j}^{\top}x+\theta_{j}$ with trainable parameters $w_{j}$, $\theta_{j}$. During training, we minimize a cost function consisting of the empirical loss as well as an additional regularization term penalizing the norms of the parameters $w_{j}$. This regularization reduces the Lipschitz bound of the quantum model w.r.t. data perturbations. We show with both theoretical and numerical results that smaller values of $\lVert w_{j}\rVert$, as encouraged by the proposed regularization, lead to more robust and generalizable quantum models.
@@
The paper is structured as follows. In Section 2, we introduce the considered class of quantum models with trainable encodings and we state their Lipschitz bound. Next, in Section 3, we use the Lipschitz bound to study robustness of quantum models and to derive a regularization strategy for robust training whose benefits are demonstrated with numerical simulations. We then derive a generalization bound which depends explicitly on the data encoding parameters and we confirm this insight numerically by showing improved generalization under the proposed regularization strategy (Section 4). Further, in Section 5, we discuss an important implication of our results on the benefits of trainable encodings for robustness and generalization. Finally, Section 6 concludes the paper. In the appendix, we provide technical proofs, details on the numerical simulations, as well as additional theoretical and numerical results.
2 Quantum models and their Lipschitz bounds
@@
with input data $x\in\mathbb{R}^{d}$, trainable parameters $\Theta_{j}=\{w_{j},\theta_{j}\}$, $w_{j}\in\mathbb{R}^{d},\theta_{j}\in\mathbb{R}$ and Hermitian generators $H_{j}$.
Depending on the choice of $H_{j}$, the operator $U_{j,\Theta_{j}}$ acts on either one or multiple qubits. The operators $U_{j,\Theta_{j}}$ give rise to the parametrized quantum circuit
@@
where $\Theta=\{\Theta_{j}\}_{j=1}^{N}$ comprises the set of trainable parameters. The quantum model considered in this paper consists of $U_{\Theta}(x)$ applied to |$\ket{0}$ and followed by a measurement w.r.t. the observable $\mathcal{M}$, i.e.,
$f_{\Theta}(x)=\braket{0}{U_{\Theta}(x)^{\dagger}\mathcal{M}U_{\Theta}(x)}{0}.$ (4)
@@
3 Robustness of quantum models
@@
Regularizing the parameters $w_{j}$ encourages small norms of the data encoding and, thereby, small values of the Lipschitz bound $L_{\Theta}$. The hyperparameter $\lambda>0$ allows for a trade off between the two objectives of a small training loss and robustness/generalization in the cost function. Note that the regularization does not involve the $\theta_{j}$’s since they do not influence the Lipschitz bound (5), an issue we discuss in more detail in Section 5. Moreover, we do not introduce an explicit dependence of the regularization on $\mathcal{M}$ since we do not optimize over the observable in this paper. We note that penalty terms similar to the proposed regularization can be used for handling hard constraints in binary optimization via the quantum approximate optimization algorithm [50, 51]. Indeed, the above regularization can be interpreted as a penalty-based relaxation of the corresponding constrained training problem, i.e., of training a quantum model with Lipschitz bound below a specific value.
@@
Figure 2: We compare robustness of quantum models trained via (9) for $\lambda\in\{0,0.2,0.5\}$ and a quantum model with fixed encoding (14). As training and test set, we draw $n=200$ and $1000$ points $x_{i}\in\mathcal{X}$, respectively, uniformly at random. To study robustness, we perturb each of the $1,000$ test data points by random noise drawn uniformly from $[-\bar{\varepsilon},+\bar{\varepsilon}]^{d}\ (d=2)$. The test accuracy in the plot is the worst case over $200$ noise samples per data point.
@@
We now evaluate our theoretical findings based on the circle classification problem from [52]: within a domain of $\mathcal{X}=[-1,+1]\times[-1,+1]$, a circle with radius $\sqrt{2/\pi}$ is drawn, and all data points inside the circle are labeled with $y=+1$, whereas points outside are labeled with $y=-1$, see Appendix C (Figure 5). For the quantum model with trainable encoding, we use general SU(2) operators and encode $w_{j}^{\top}x+\theta_{j}$ into the first two rotation angles. We repeat this encoding for each of the considered $3$ qubits, followed by nearest neighbour entangling gates based on CNOTs. Such a layer is then repeated $3$ times. As observable, we use $\mathcal{M}=Z\otimes Z\otimes Z$. The resulting circuit is illustrated in Appendix C (Figure 7). As norms in the regularized training problem (9), we employ $2$-norms, but we note that exploring different choices is an interesting future direction. For example, a $1$-norm regularization may enforce sparsity and, thereby, simplify implementations on NISQ hardware.
@@
The numerical results for the robustness simulations are shown in Figure 2, where we compare the worst-case test accuracy and Lipschitz bound of three trained models with different regularization parameter $\lambda\in\{0,0.2,0.5\}$. Additionally, the plot shows the accuracy of a quantum model with fixed encoding for the same numbers of qubits and layers (see (14) and Appendix C for details). The worst-case test accuracy of all models is obtained by sampling different noise samples $\varepsilon$ from $[-\bar{\varepsilon},+\bar{\varepsilon}]$ and, thereby, it approximates the accuracy resulting from an adversarial attack norm-bounded by $\bar{\varepsilon}$. As expected, all four models deteriorate with increasing noise level. For zero noise level $\bar{\varepsilon}=0$, the model with the largest regularization parameter $\lambda=0.5$ (and, hence, the smallest Lipschitz bound $L_{\Theta}=3.67$) has a smaller test accuracy than the non-regularized model with $\lambda=0$. This can be explained by a decrease in the training accuracy that is caused by the additional regularization in the cost. For increasing noise levels, however, the enhanced robustness outweighs the loss of training performance and, therefore, the model with $\lambda=0.5$ outperforms the model with $\lambda=0$. The fixed-encoding model achieves comparable performance to the trainable-encoding model with $\lambda=0.5$ for small noise and the worst performance among all models for high noise. These observations can be explained by the reduced expressivity but high Lipschitz bound of the fixed-encoding model. Finally, the model with $\lambda=0.2$ almost always outperforms the model with $\lambda=0$ and, in particular, it yields a higher test accuracy for small noise levels. This can be explained by the improved generalization performance caused by the regularization, an effect we discuss in more detail in the following.
4 Generalization of quantum models
@@
The detailed version and proof of Theorem 4.1 are provided in Appendix B. Generalization bounds as in (10) quantify the ability of $f_{\Theta}$ to generalize beyond the available data. The bound (10) depends on the data encoding via $\sum_{j=1}^{N}\lVert w_{j}\rVert\lVert H_{j}\rVert$ and on the observable via $\lVert\mathcal{M}\rVert$. In particular, $f_{\Theta}$ achieves a small generalization error if its Lipschitz bound $L_{\Theta}$ is small and the size $n$ of the data set is large. Note, however, the following fundamental trade-off: A too small Lipschitz bound $L_{\Theta}$ may limit the expressivity of $f_{\Theta}$ and, therefore, lead to a high empirical risk $\mathcal{R}_{n}(f_{\Theta})$, in which case the generalization bound (10) is meaningless. In conclusion, Theorem 4.1 implies a small expected risk $\mathcal{R}(f_{\Theta})$ if $n$ is large, $L_{\Theta}$ is small, and $f_{\Theta}$ has a small empirical risk $\mathcal{R}_{n}(f_{\Theta})$. In contrast to existing generalization bounds [24, 45, 46, 47, 48], the bound (10) is not uniform and explicitly involves the Lipschitz bound (5), i.e., the parameters of the data encoding. Hence, Theorem 4.1 does not share the limitations of uniform QML generalization bounds [49].
We evaluate the generalization performance of the trainable encoding again on the circle classification problem. The training setup is identical to the robustness simulations and the numerical results are shown in Figure 3. Increasing the regularization parameter $\lambda$ decreases the Lipschitz bound $L_{\Theta}$ of the trained model. In accordance with the generalization bound (10), this reduction of $L_{\Theta}$ improves generalization performance with the maximum test accuracy at $\lambda=0.15$. Beyond this value, the regularization causes a too small Lipschitz bound, limiting expressivity and, therefore, decreasing the training accuracy. As a result, the test accuracy decreases as well. This illustrates the role of $\lambda$ as a hyperparameter: Regularization does not always improve performance, but there is a sweet spot for $\lambda$ at which both superior generalization and robustness over the unregularized setup (i.e., $\lambda=0$) can be obtained. In practice, the hyperparameter $\lambda$ can be tuned, e.g., via cross-validation.
5 Benefits of trainable encodings
@@
$f^{\mathrm{f}}_{\phi}(x)=\braket{0}{U^{\mathrm{f}}_{\phi}(x)^{\dagger}\mathcal{M}U^{\mathrm{f}}_{\phi}(x)}{0},$ (14)
@@ A.1 Simple Lipschitz bound based on concatenation
$g_{\mathrm{meas}}(z_{\mathrm{m}})$ $=\braket{z_{\mathrm{m}}}{\mathcal{M}}{z_{\mathrm{m}}},$ (19)
$g_{\mathrm{unitary}}(z_{\mathrm{u}})$ $=e^{-iz_{\mathrm{u},N}H_{N}}\cdots e^{-iz_{\mathrm{u},1}H_{1}}\ket{0},$ (20)
@@
$\frac{\mathrm{d}g_{\mathrm{meas}}(z_{\mathrm{m}})}{\mathrm{d}z_{\mathrm{m}}}=2\bra{z_{\mathrm{m}}}\mathcal{M}.$ (24)
@@ A.2 Proof that (5) is a Lipschitz bound
We have thus shown that the Jacobian $J_{\Theta}(x)$ of the map $U_{\Theta}(x)\ket{0}$ is given by
@@
$=$ $-i\left(H_{N}U_{\Theta}(x)\ket{0}w_{N}^{\top}+\dots+U_{\Theta}(x)H_{1}\ket{0}w_{1}^{\top}\right).$
Using that |$\ket{0}$ has unit norm, that the $U_{j,\Theta_{j}}$’s are unitary, as well as the triangle inequality, the norm of $J_{\Theta}(x)$ is bounded as
@@
Thus, $\sum_{j=1}^{N}\lVert w_{j}\rVert\lVert H_{j}\rVert$ is a Lipschitz bound of $U_{\Theta}(x)\ket{0}$ [59, p. 356].
Finally, $f_{\Theta}(x)$ is a concatenation of $U_{\Theta}(x)\ket{0}$ and the function $z\mapsto\braket{z}{\mathcal{M}}{z}$, which admits the Lipschitz bound $2\lVert\mathcal{M}\rVert$, compare (25). Hence, a Lipschitz bound of $f_{\Theta}$ can be obtained as the product of these two individual bounds, i.e., as in (5).
Appendix B Full version and proof of Theorem 4.1
@@
Let us now combine the Lipschitz bound (5) and Lemma B.1 to state a tailored generalization bound for the considered class of quantum models $f_{\Theta}$, thus proving Theorem 4.1.
@@
In contrast to existing QML generalization bounds [24, 45, 46, 47, 48], Theorem B.1 explicitly highlights the role of the model parameters via the Lipschitz bound. Using the additional flexibility of the parameter $\gamma$, it can be shown that the generalization bound (40) converges to zero when the data length $n$ approaches infinity. To this end, we use [60, Lemma 6.27] to upper bound the covering number

### v4

@@ Abstract
Adversarial robustness and generalization are both crucial properties of reliable machine learning models. In this paper, we study these properties in the context of quantum machine learning based on Lipschitz bounds. We derive parameter-dependent Lipschitz bounds for quantum models with trainable encoding, showing that the norm of the data encoding has a crucial impact on the robustness against data perturbations. Further, we derive a bound on the generalization error which explicitly involves the parameters of the data encoding. Based on these theoretical results, we propose a practical strategy for training robust and generalizable quantum models by regularizing the Lipschitz bound in the cost. Moreover, we show that, for fixed and non-trainable encodings, as those frequently employed in quantum machine learning, the Lipschitz bound cannot be influenced by tuning the parameters. Thus, trainable encodings are crucial for systematically adapting robustness and generalization during training. The practical implications of our theoretical findings are illustrated with numerical results.
preprint: APS/123-QED
I Introduction
@@ A Lipschitz bound of a model $f$ is any $L>0$ satisfying
for all $x_{1},x_{2}\in\mathcal{D}\subseteq\mathbb{R}^{d}$, where $d$ is the data dimension. By definition, Lipschitz bounds quantify the worst-case output change that can be caused by data perturbations and, thus, they provide a useful measure of adversarial robustness. Therefore, they are a well-established tool for characterizing robustness and generalization properties of ML models [2, 6, 9, 10, 11, 12, 13, 14, 15]. Lipschitz bounds cannot only be used to better understand these two properties, but they also allow one to improve them by regularizing the Lipschitz bound during training [2, 6, 16, 17, 18].
Figure 1: Schematic illustration of the quantum model and training setup considered in this work for an exemplary Fashion MNIST data set [19]. The data, $x$, enter the quantum circuit via a trainable encoding, i.e., they are encoded into unitary operators $U_{j,\Theta_{j}}(x)$ via an affine function $w_{j}^{\top}x+\theta_{j}$ with trainable parameters $w_{j}$, $\theta_{j}$. During training, we minimize a cost function consisting of the empirical loss as well as an additional regularization term penalizing the norms of the parameters $w_{j}$. This regularization reduces the Lipschitz bound of the quantum model w.r.t. data perturbations and, thereby, encourages improved robustness and generalization properties.
@@
The paper is structured as follows. In Section II, we introduce the considered class of quantum models with trainable encodings and we state their Lipschitz bound. Next, in Section III, we use the Lipschitz bound to study robustness of quantum models and to derive a regularization strategy for robust training whose benefits are demonstrated with numerical simulations. We then derive a generalization bound which depends explicitly on the data encoding parameters and we confirm this insight numerically by showing improved generalization under the proposed regularization strategy (Section IV). Further, in Section V, we discuss an important implication of our results on the benefits of trainable encodings for robustness and generalization. Finally, Section VI concludes the paper. In the appendix, we provide technical proofs, details on the numerical simulations, as well as additional theoretical and numerical results.
II Quantum models and their Lipschitz bounds
@@
with input data $x\in\mathbb{R}^{d}$, trainable parameters $\Theta_{j}=\{w_{j},\theta_{j}\}$, $w_{j}\in\mathbb{R}^{d},\theta_{j}\in\mathbb{R}$ and fixed Hermitian generators $H_{j}$. The generators $H_{j}$ are user-chosen, see [43] for references with guidelines. Depending on the choice of $H_{j}$, the operator $U_{j,\Theta_{j}}$ acts on either one or multiple qubits. The operators $U_{j,\Theta_{j}}$ give rise to the parametrized quantum circuit
@@
[…]
$f_{\Theta}(x)=\braket{0|U_{\Theta}(x)^{\dagger}\mathcal{M}U_{\Theta}(x)|0}.$ (4)
@@
III Robustness of quantum models
@@
Regularizing the parameters $w_{j}$ encourages small norms of the data encoding and, thereby, small values of the Lipschitz bound $L_{\Theta}$. We weight the parameter norms $\lVert w_{j}\rVert$ by $\lVert H_{j}\rVert$ due to their joint occurrence in (5). The hyperparameter $\lambda>0$ allows for a trade off between the two objectives of a small training loss and robustness/generalization in the cost function. Note that the regularization does not involve the $\theta_{j}$’s since they do not influence the Lipschitz bound (5), an issue we discuss in more detail in Section V. Moreover, we do not introduce an explicit dependence of the regularization on $\mathcal{M}$ since we do not optimize over the observable in this paper. We note that penalty terms similar to the proposed regularization can be used for handling hard constraints in binary optimization via the quantum approximate optimization algorithm [50, 51]. Indeed, the above regularization can be interpreted as a penalty-based relaxation of the corresponding constrained training problem, i.e., of training a quantum model with Lipschitz bound below a specific value.
@@
Figure 2: We compare robustness of quantum models trained via (9) for $\lambda\in\{0,0.2,0.5\}$ and a quantum model with fixed encoding (14). As training and test set, we draw $n=200$ and $1000$ points $x_{i}\in\mathcal{X}$, respectively, uniformly at random. To study robustness, we perturb each of the $1,000$ test data points by random noise drawn uniformly from $\bar{\varepsilon},+\bar{\varepsilon}]^{d}\ (d=2)$. The test accuracy in the plot is the worst case over $200$ noise samples per data point.
@@
We now evaluate our theoretical findings based on the circle classification problem from [52]: within a domain of $\mathcal{X}=[-1,+1]\times[-1,+1]$, a circle with radius $\sqrt{2/\pi}$ is drawn, and all data points inside the circle are labeled with $y=+1$, whereas points outside are labeled with y=1y=-1, see Appendix C (Figure 5). For the quantum model with trainable encoding, we use general SU(2) operators and encode $w_{j}^{\top}x+\theta_{j}$ into the first two rotation angles. We repeat this encoding for each of the considered $3$ qubits, followed by nearest neighbour entangling gates based on CNOTs. Such a layer is then repeated $3$ times. As observable, we use $\mathcal{M}=Z\otimes Z\otimes Z$. The resulting circuit is illustrated in Appendix C (Figure 7). As norms in the regularized training problem (9), we employ $2$-norms, but we note that exploring different choices is an interesting future direction. For example, regularization with non-squared norms (e.g., a $1$-norm) may enforce sparsity of the trained QML model and can, thereby, simplify its implementation on NISQ hardware.
@@
The numerical results for the robustness simulations are shown in Figure 2, where we compare the worst-case test accuracy and Lipschitz bound of three trained models with different regularization parameter $\lambda\in\{0,0.2,0.5\}$. Additionally, the plot shows the accuracy of a trained quantum model with fixed encoding for the same numbers of qubits and layers (see (14) and Appendix C for details). The worst-case test accuracy of all models is obtained by sampling different noise samples $\varepsilon$ from $\bar{\varepsilon},+\bar{\varepsilon}]$. This procedure amounts to finding adversarial noise samples and, therefore, the resulting worst-case test accuracy (approximately) quantifies the adversarial robustness against attacks which are norm-bounded by $\bar{\varepsilon}$. As expected, all four models deteriorate with increasing noise level. For zero noise level $\bar{\varepsilon}=0$, the model with the largest regularization parameter $\lambda=0.5$ (and, hence, the smallest Lipschitz bound $L_{\Theta}=3.67$) has a smaller test accuracy than the non-regularized model with $\lambda=0$. This can be explained by a decrease in the training accuracy that is caused by the additional regularization in the cost. For increasing noise levels, however, the enhanced robustness outweighs the loss of training performance and, therefore, the model with $\lambda=0.5$ outperforms the model with $\lambda=0$. The fixed-encoding model achieves comparable performance to the trainable-encoding model with $\lambda=0.5$ for small noise and the worst performance among all models for high noise. These observations can be explained by the high Lipschitz bound of the fixed-encoding model as well as its reduced expressivity, i.e., its limited ability to approximate functions from data due to the fixed encoding parameters $w_{j}$. Finally, the model with $\lambda=0.2$ almost always outperforms the model with $\lambda=0$ and, in particular, it yields a higher test accuracy for small noise levels. This can be explained by the improved generalization performance caused by the regularization, an effect we discuss in more detail in the following.
IV Generalization of quantum models
@@
The detailed version and proof of Theorem IV.1 are provided in Appendix B. Generalization bounds as in (10) quantify the ability of $f_{\Theta}$ to generalize beyond the available data. The bound (10) depends on the data encoding via $\sum_{j=1}^{N}\lVert w_{j}\rVert\lVert H_{j}\rVert$ and on the observable via $\lVert\mathcal{M}\rVert$. In particular, $f_{\Theta}$ achieves a small generalization error if its Lipschitz bound $L_{\Theta}$ is small and the size $n$ of the data set is large. Note, however, the following fundamental trade-off: A too small Lipschitz bound $L_{\Theta}$ may limit the expressivity of $f_{\Theta}$ and, therefore, lead to a high empirical risk $\mathcal{R}_{n}(f_{\Theta})$, in which case the generalization bound (10) is meaningless. In conclusion, Theorem IV.1 implies a small expected risk $\mathcal{R}(f_{\Theta})$ if $n$ is large, $L_{\Theta}$ is small, and $f_{\Theta}$ has a small empirical risk $\mathcal{R}_{n}(f_{\Theta})$. In contrast to existing generalization bounds [24, 45, 46, 47, 48], the bound (10) is not uniform and explicitly involves the Lipschitz bound (5), i.e., the parameters of the data encoding. Hence, Theorem IV.1 does not share the limitations of uniform QML generalization bounds [49] and it can be used to systematically influence the generalization performance during training via regularization. In particular, according to Theorem IV.1, the regularized training problem (9) encourages models with improved generalization properties, where the hyperparameter $\lambda$ trades off between the empirical risk $\mathcal{R}_{n}(f_{\Theta})$ and the generalization bound (10). Extending our results by studying the direct impact of $\lambda$ on the generalization performance is an interesting next step, which we expect to be non-trivial due to the non-convexity of the loss function in (9), compare [53]. In practice, the hyperparameter $\lambda$ can be tuned, e.g., via cross-validation. We discuss and interpret the impact of the hyperparameter $\lambda$ in more detail with the following numerical results.
We evaluate the generalization performance of the trainable encoding again on the circle classification problem. The training setup is identical to the robustness simulations and the numerical results are shown in Figure 3. Increasing the regularization parameter $\lambda$ decreases the Lipschitz bound $L_{\Theta}$ of the trained model. In accordance with the generalization bound (10), this reduction of $L_{\Theta}$ improves generalization performance with the maximum test accuracy at $\lambda=0.15$. Beyond this value, the regularization causes a too small Lipschitz bound, limiting expressivity and, therefore, decreasing the training accuracy. As a result, the test accuracy decreases as well. This illustrates the role of $\lambda$ as a hyperparameter: Regularization does not always improve performance, but there is a sweet spot for $\lambda$ at which both superior generalization and robustness over the unregularized setup (i.e., $\lambda=0$) can be obtained.
V Benefits of trainable encodings
@@
$f^{\mathrm{f}}_{\phi}(x)=\braket{0|U^{\mathrm{f}}_{\phi}(x)^{\dagger}\mathcal{M}U^{\mathrm{f}}_{\phi}(x)|0},$ (14)
@@ A.1 Simple Lipschitz bound based on concatenation
$g_{\mathrm{meas}}(z_{\mathrm{m}})$ $=\braket{z_{\mathrm{m}}|\mathcal{M}|z_{\mathrm{m}}},$ (19)
[…]
@@
[…]
@@ A.2 Proof that (5) is a Lipschitz bound
[…]
@@
[…]
@@
[…]
Appendix B Full version and proof of Theorem IV.1
@@
Let us now combine the Lipschitz bound (5) and Lemma B.1 to state a tailored generalization bound for the considered class of quantum models $f_{\Theta}$, thus proving Theorem IV.1.
@@
In contrast to existing QML generalization bounds [24, 45, 46, 47, 48], Theorem B.1 explicitly highlights the role of the model parameters via the Lipschitz bound. Using the additional flexibility of the parameter $\gamma$, it can be shown that the generalization bound (40) converges to zero when the data length $n$ approaches infinity. To this end, we use [61, Lemma 6.27] to upper bound the covering number